A woman is now 4 times older than her daughter. Six years ago, the product of their ages was 136. Find their present ages.
Answer by Merciful(5)

Let x = present daughter's age
Then 4x = woman's present age
x - 6 = daughter's age six years ago
4x - 6 = woman's age six years ago
Therefore
( x - 6 )( 4x - 6 ) = 136
4x^2 - 6x - 24x + 36 = 136
4x^2 - 30x - 100 = 0
2x^2 - 15x - 50 = 0
( 2x + 5 )( x - 10 ) = 0
x = 10 or x = -5/2 N/A the age can't be negative
Daughter's age = 10
And woman's age = 4 X 10 = 40